Formula for getting time of sunrise at a particular location?

Grade:9

1 Answers

SAHIL
7 years ago
FOR THAT YOU CAN USE SUNRISE EQUATIO BUT YOU HAVE TO CONSIDER THAT NOON IS EXACTLY HALF WAY BETWEEN  SUNRISE AND SUNSET
\cos \omega _{\circ }=-\tan \phi \times \tan \delta

where:

 is the sun DECLINATION\delta {\displaystyle \delta }  value is taken);positive value is taken) or sunset (when negative is the hour angle  at either sunrise (when \omega _{\circ }{\displaystyle \omega _{\circ }}